OpenStudy (anonymous):

Find the sum. 7x + 15x = 21x 21x2 22x 22x2

OpenStudy (abdullahm):

Hi

OpenStudy (abdullahm):

What is 7+15 = ?

OpenStudy (dmndlife24):

add like terms bruh

OpenStudy (anonymous):

22

OpenStudy (abdullahm):

So, do you know the distribution property?

OpenStudy (anonymous):

no

OpenStudy (abdullahm):

ok, it's \(\sf a(b+c) = ab+ac\)

OpenStudy (abdullahm):

So for your question we can use the distributive property and simplify it. \(\sf 7x + 15x\) What thing is common in between 7x and 15x ?

OpenStudy (abdullahm):

Hint: It's the variable.

OpenStudy (abdullahm):

@jpezzyy ??

OpenStudy (anonymous):

22x